OpenStudy (anonymous):

A package of spaghetti weighs 266 grams. How many ounces does the package weigh? 1 oz approximately symbol 28 g (3 points)

OpenStudy (anonymous):

\[\frac{1oz}{28g}=\frac{??oz}{266g}\]

OpenStudy (anonymous):

1 oz = 28 g 1/28 oz = 1 g Keep going until you get 266 g = ??? oz

OpenStudy (anonymous):

mmm k